If a firm's demand function is P = 100 - q, what quantities need to be sold to bring in a total revenue of £1000


1
Expert's answer
2021-11-02T10:43:07-0400
TR=pq=(100q)qTR=pq=(100-q)q

Substitute


1000=100qq2,q01000=100q-q^2, q\geq0

q2100q+1000=0q^2-100q+1000=0

D=(100)24(1000)=6000D=(-100)^2-4(1000)=6000

q=100±60002q=\dfrac{100\pm\sqrt{6000}}{2}

q=50±1015q=50\pm10\sqrt{15}


q111,q289q_1\approx11, q_2\approx89